OpenStudy (anonymous):

g9x0=x suqared + 13 , g(-3) HELP !!!!!!!!!

OpenStudy (anonymous):

I dont know what that is? \[g(x) = x^2 + 13 \] ?? if so then \[g(-3) = (-3)^2 + 13 = 9 + 13 = 22\]